OpenStudy (anonymous):

anybody can help me with Q5 of midterm....

OpenStudy (anonymous):

for Q5: Put the given value in this: 1)\[(Vs - Vout)^{2} = (K*R)^{2} * (Vin - 2*(Vs-Vout))\] 2) \[(Vin / Vout) = -2 - (2/((K*R)^{2})) *(Vs - Vout) \] for Second equation: after finding the Vin/Vout, just take the reciprocal of it because in the answer they required Vout/Vin

OpenStudy (anonymous):

thankss...... and you have any idea of Vout of Q4.... i dont know how i can put iRl for node eq